# DSR-Bench: Evaluating the Structural Reasoning Abilities of LLMs via Data Structures

### v1

@@ Appendix A Appendix
A.1 Additional related works on LLM benchmarking
Large language models (LLMs) have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations. Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding [49, 9], text generation [33, 41], reasoning [9, 46], and machine translation [53]. Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness [42, 31, 19, 24, 57, 29]. Specialized benchmarks have also emerged in scientific and technical domains, including mathematics [50, 25], programming [50, 16, 3, 58], data analysis [43, 27, 50, 44, 17], and medicine [23, 26, 30, 52, 18].
While several studies examine how LLMs convert unstructured inputs into tabular or relational formats [44, 17], our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs? To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
@@ A.2 Details of data structures and operations
In this section, we list the data structures and the corresponding operations tested in Table 8. We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
@@
| LRU Cache | Least-recently-used | Cache (Evict, Add) | Web broswers | |
@@
| Associative | Hashmap | Key-value storage | Compound | Large-scaled storage |
@@
| K-D Heap* | Hierarchical, complete binary tree, high dimensional priority | Compound | GPU job scheduling | |
@@ A.3 Examples of prompting strategies
This method expliclty adds a step attribute in the structured output, guiding the model to produce operations in a sequential and interpretable manner.
@@ A.4 Accuracy by task and length level across all models
In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level. Table 9 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels. Table 10 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite. For detailed per-model results across all tasks and length levels, see Table 11 (o4-mini), Table 12 (Gemini-2.5-Pro), Table 13 (Claude-3.7-Sonnet), Table 14 (DeepSeek-R1), Table 15 (GPT-4.1), Table 16 (Gemini-2.0-Flash), Table 17 (Claude-3.5-Sonnet), Table 18 (DeepSeek-V3), and Table 19 (Llama-3.3).
@@
| Category | Data Structure | Length | o4-mini | Gemini-2.5-Pro | Claude-3.7-Sonnet | DeepSeek-R1 |
| Temporal | Priority Queue | Short | 0.89 | 0.87 | 0.70 | 0.92 |
| Medium | 0.47 | 0.43 | 0.11 | 0.54 | ||
| Long | 0.30 | 0.24 | 0.04 | 0.48 | ||
| Associative | Trie | Short | 0.99 | 1.00 | 0.23 | 0.93 |
| Medium | 0.73 | 0.82 | 0.00 | 0.50 | ||
| Long | 0.32 | 0.59 | 0.00 | 0.05 | ||
| Suffix Tree | Short | 0.96 | 0.74 | 0.96 | 1.00 | |
| Medium | 0.87 | 0.92 | 0.98 | 0.98 | ||
| Long | 0.37 | 0.89 | 0.79 | 0.90 | ||
| Skip List | Short | 0.84 | 0.93 | 0.87 | 0.89 | |
| Medium | 0.60 | 0.64 | 0.76 | 0.63 | ||
| Long | 0.41 | 0.61 | 0.61 | 0.54 | ||
| Hierarchical | Heap | Short | 0.61 | 0.67 | 0.70 | 0.63 |
| Medium | 0.72 | 0.55 | 0.37 | 0.53 | ||
| Long | 0.71 | 0.62 | 0.13 | 0.27 | ||
| Red Black Tree | Short | 0.92 | 0.97 | 0.69 | 0.86 | |
| Medium | 0.67 | 0.73 | 0.11 | 0.63 | ||
| Long | 0.37 | 0.60 | 0.12 | 0.37 | ||
| B+ Tree | Short | 0.99 | 0.97 | 0.70 | 0.49 | |
| Medium | 0.98 | 0.93 | 0.32 | 0.23 | ||
| Long | 0.94 | 0.96 | 0.13 | 0.21 | ||
| K-D Heap | Short | 0.22 | 0.07 | 0.11 | 0.23 | |
| Medium | 0.09 | 0.08 | 0.03 | 0.08 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.01 | ||
| K-D Tree | Short | 0.59 | 0.64 | 0.00 | 1.00 | |
| Medium | 0.43 | 0.81 | 0.00 | 0.34 | ||
| Long | 0.38 | 0.64 | 0.00 | 0.01 | ||
| Network | DSU | Short | 1.00 | 1.00 | 1.00 | 1.00 |
| Medium | 1.00 | 1.00 | 1.00 | 1.00 | ||
| Long | 0.94 | 0.99 | 0.98 | 0.99 | ||
| Hybrid | Bloom Filter | Short | 1.00 | 0.99 | 0.44 | 0.99 |
| Medium | 0.77 | 1.00 | 0.03 | 0.92 | ||
| Long | 0.07 | 0.69 | 0.00 | 0.31 | ||
| DAWG | Short | 0.49 | 0.07 | 0.17 | 0.40 | |
| Medium | 0.20 | 0.14 | 0.00 | 0.02 | ||
| Long | 0.06 | 0.00 | 0.00 | 0.00 | ||
| Geom Graph | Short | 0.37 | 0.30 | 0.07 | 0.36 | |
| Medium | 0.10 | 0.10 | 0.00 | 0.01 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.01 |
A.4.1 Performance of o4-mini
@@
| Geom Graph* | Construct | 0.37 (0.00) | 0.10 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 1.00 (0.00) | 0.77 (0.09) | 0.07 (0.00) |
@@
| Priority Queue* | Compound | 0.87 (0.03) | 0.43 (0.03) | 0.24 (0.05) | |
| Associative | Hashmap | Compound | 0.90 (0.06) | 0.51 (0.04) | 0.50 (0.07) |
| Trie* | Compound | 1.00 (0.00) | 0.82 (0.05) | 0.59 (0.08) | |
| Suffix Tree* | Construct | 0.74 (0.08) | 0.92 (0.08) | 0.89 (0.02) | |
| Skip List* | Compound | 0.93 (0.03) | 0.64 (0.05) | 0.61 (0.02) | |
| Associative | BST | Insert | 0.99 (0.02) | 0.94 (0.05) | 0.96 (0.02) |
| Remove | 0.99 (0.02) | 0.84 (0.05) | 0.90 (0.03) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
@@
| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Depth | 0.22 (0.02) | 0.17 (0.07) | 0.18 (0.04) | ||
| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | ||
| Heap* | Compound | 0.91 (0.04) | 0.72 (0.11) | 0.71 (0.05) | |
| Heapify | 0.43 (0.21) | 0.37 (0.03) | 0.53 (0.09) | ||
| RB Tree* | Construct | 0.83 (0.05) | 0.60 (0.03) | 0.28 (0.07) | |
| Compound | 0.91 (0.02) | 0.47 (0.06) | 0.18 (0.08) | ||
| B+ Tree* | Compound | 0.97 (0.06) | 0.93 (0.03) | 0.96 (0.05) | |
| K-D Tree* | Construct | 0.64 (0.11) | 0.81 (0.05) | 0.67 (0.12) | |
| K-D Heap* | Compound | 0.07 (0.00) | 0.08 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.98 (0.04) | 0.99 (0.02) | 0.86 (0.12) |
| Depth-First Traversal | 0.98 (0.02) | 0.94 (0.04) | 0.86 (0.05) | ||
| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 0.99 (0.02) | |
| Geom Graph* | Construct | 0.30 (0.03) | 0.10 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.99 (0.02) | 1.00 (0.00) | 0.69 (0.02) |
| DAWG* | Compound | 0.07 (0.02) | 0.14 (0.08) | 0.00 (0.00) |
A.4.3 Performance of Claude-3.7-Sonnet
@@
| Geom Graph* | Construct | 0.07 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) |
@@ A.4.4 Performance of DeepSeek-R1
| Geom Graph* | Construct | 0.36 (0.02) | 0.01 (0.02) | 0.01 (0.02) | |
| Network | Bloom Filter* | Compound | 0.99 (0.02) | 0.92 (0.02) | 0.31 (0.02) |
@@
| Geom Graph* | Construct | 0.20 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
A.4.6 Performance of Gemini-2.0-Flash
@@
| Geom Graph* | Construct | 0.06 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
A.4.7 Performance of Claude-3.5-Sonnet
@@
| Geom Graph* | Construct | 0.06 (0.04) | 0.00 (0.00) | 0.00 (0.00) | |
| Network | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |

### v2

@@ Appendix A Appendix
A.1 Additional related work on LLM benchmarking
Large language models (LLMs) have demonstrated remarkable performance across a wide range of applications, prompting growing interest in understanding their capabilities and limitations. Recent efforts have focused on systematically benchmarking LLMs on core natural language processing tasks, including language understanding (Wang et al., 2024; Feng et al., 2024), text generation (Maynez et al., 2023; Singh et al., 2024), reasoning (Feng et al., 2024; Valmeekam et al., 2023), and machine translation (Yao et al., 2023). Additionally, some benchmarks evaluate ethical dimensions such as robustness, bias, and trustworthiness (Siska et al., 2024; Luo et al., 2024; Koo et al., 2023; Li et al., 2024b; Zhang et al., 2024b; Liu et al., 2023c). Specialized benchmarks have also emerged in scientific and technical domains, including mathematics (White et al., 2025; Liu et al., 2024a), programming (White et al., 2025; Jain et al., 2025; Chen et al., 2021; Zheng et al., 2023), data analysis (Sui et al., 2024; Liu et al., 2023b; White et al., 2025; Tang et al., 2023; Jiang et al., 2023), and medicine (Li et al., 2024a; Liu et al., 2023a; Longwell et al., 2024; Xu et al., 2024b; Kanithi et al., 2024).
While several studies examine how LLMs convert unstructured inputs into tabular or relational formats (Tang et al., 2023; Jiang et al., 2023), our work explores a distinct question: How well can LLMs construct, manipulate, and reason about classic data structures such as stacks, trees, and graphs? To our knowledge, this is the first comprehensive benchmark targeting this capability, offering a conceptually different evaluation from prior work on structured data.
@@ A.2 Details of data structures and operations
In this section, we list the data structures and the corresponding operations tested in Table˜10. We then provide detailed descriptions of each data structure and explain how we specify their implementations to eliminate ambiguity.
@@
| LRU Cache | Least-recently-used | Cache (Evict, Add) | Web browsers | |
@@
| Associative | Hashmap | Key-value storage | Compound | Large-scale storage |
@@
| K-D Heap* | Hierarchical, complete binary tree, high-dimensional priority | Compound | GPU job scheduling | |
@@ A.3 Examples of prompting strategies
This method explicitly adds a steps attribute in the JSON schema of Structured Output, guiding the model to produce operations in a sequential and interpretable manner. Below is an example schema for an array task:
class Step(BaseModel): explanation: str output: str class ArraySchema(BaseModel): steps: list[Step] final_answer: int
@@ A.4 Accuracy by task and length level across all models
In this section, we provide supplementary accuracy tables for all models in DSR-Bench, broken down by task and length level. Table˜11 summarizes the accuracy of instruction-tuned models on a subset of basic data structures across different length levels. Table˜12 presents the accuracy of reasoning models on selected data structures from the DSR-Bench-challenge suite. For detailed per-model results across all tasks and length levels, see Table˜13 (GPT-5), Table˜14 (o4-mini), Table˜15 (Gemini-2.5-Pro), Table˜16 (Claude-3.7-Sonnet), Table˜17 (DeepSeek-R1), Table˜18 (GPT-4.1), Table˜19 (Gemini-2.0-Flash), Table˜20 (Claude-3.5-Sonnet), Table˜21 (DeepSeek-V3), and Table˜22 (Llama-3.3).
@@
| Category | Data Structure | Length | o4-mini | Gemini-2.5 Pro | Claude-3.7 Sonnet | Deepseek R1 | GPT-5 (med) |
| Temporal | Priority Queue | Short | 0.89 | 0.89 | 0.70 | 0.92 | 0.84 |
| Medium | 0.47 | 0.41 | 0.11 | 0.54 | 0.44 | ||
| Long | 0.30 | 0.23 | 0.04 | 0.48 | 0.28 | ||
| Associative | Trie | Short | 0.99 | 0.93 | 0.23 | 0.93 | 0.97 |
| Medium | 0.73 | 0.56 | 0.00 | 0.50 | 0.94 | ||
| Long | 0.32 | 0.37 | 0.00 | 0.05 | 0.92 | ||
| Suffix Tree | Short | 0.96 | 0.96 | 0.96 | 1.00 | 1.00 | |
| Medium | 0.87 | 0.89 | 0.98 | 0.98 | 1.00 | ||
| Long | 0.37 | 0.86 | 0.79 | 0.90 | 0.95 | ||
| Skip List | Short | 0.84 | 0.94 | 0.87 | 0.89 | 0.88 | |
| Medium | 0.60 | 0.87 | 0.76 | 0.63 | 0.66 | ||
| Long | 0.41 | 0.53 | 0.61 | 0.54 | 0.51 | ||
| Hierarchical | Heap | Short | 0.61 | 0.57 | 0.70 | 0.63 | 0.78 |
| Medium | 0.72 | 0.28 | 0.37 | 0.53 | 0.71 | ||
| Long | 0.71 | 0.29 | 0.13 | 0.27 | 0.71 | ||
| Red Black Tree | Short | 0.92 | 0.91 | 0.69 | 0.86 | 0.92 | |
| Medium | 0.67 | 0.47 | 0.11 | 0.63 | 0.79 | ||
| Long | 0.37 | 0.08 | 0.12 | 0.37 | 0.59 | ||
| B+ Tree | Short | 0.99 | 1.00 | 0.70 | 0.49 | 0.97 | |
| Medium | 0.98 | 0.97 | 0.32 | 0.23 | 0.99 | ||
| Long | 0.94 | 0.94 | 0.13 | 0.21 | 0.98 | ||
| K-D Heap | Short | 0.22 | 0.23 | 0.11 | 0.23 | 0.23 | |
| Medium | 0.09 | 0.06 | 0.03 | 0.08 | 0.08 | ||
| Long | 0.00 | 0.00 | 0.00 | 0.01 | 0.00 | ||
| K-D Tree | Short | 0.59 | 0.96 | 0.00 | 1.00 | 0.97 | |
| Medium | 0.43 | 0.64 | 0.00 | 0.34 | 0.90 | ||
| Long | 0.38 | 0.16 | 0.00 | 0.01 | 0.67 | ||
| Network | DSU | Short | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 |
| Medium | 1.00 | 0.99 | 1.00 | 1.00 | 1.00 | ||
| Long | 0.94 | 0.90 | 0.98 | 0.99 | 1.00 | ||
| Hybrid | Bloom Filter | Short | 1.00 | 0.94 | 0.44 | 0.99 | 1.00 |
| Medium | 0.77 | 0.97 | 0.03 | 0.92 | 0.84 | ||
| Long | 0.07 | 0.66 | 0.00 | 0.31 | 0.47 | ||
| DAWG | Short | 0.49 | 0.61 | 0.17 | 0.40 | 1.00 | |
| Medium | 0.20 | 0.08 | 0.00 | 0.02 | 0.03 | ||
| Long | 0.06 | 0.01 | 0.00 | 0.00 | 0.00 | ||
| Geom Graph | Short | 0.37 | 0.19 | 0.07 | 0.36 | 0.21 | |
| Medium | 0.10 | 0.17 | 0.00 | 0.01 | 0.08 | ||
| Long | 0.00 | 0.02 | 0.00 | 0.01 | 0.19 |
A.4.1 Performance of GPT-5
| Category | Data Structure | Operation | Short | Medium | Long |
| Linear | Array | Access | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Delete | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Reverse | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Search | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Temporal | Stack | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Queue | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| LRU Cache | Cache | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| Priority Queue* | Compound | 0.84 (0.02) | 0.44 (0.02) | 0.28 (0.07) | |
| Associative | Hashmap | Compound | 1.00 (0.00) | 0.89 (0.02) | 0.71 (0.10) |
| Trie* | Compound | 0.97 (0.03) | 0.94 (0.05) | 0.92 (0.07) | |
| Suffix Tree* | Construct | 1.00 (0.00) | 1.00 (0.00) | 0.95 (0.02) | |
| Skip List* | Compound | 0.88 (0.05) | 0.66 (0.08) | 0.51 (0.05) | |
| Hierarchical | BST | Insert | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) |
| Remove | 0.98 (0.02) | 1.00 (0.00) | 0.98 (0.02) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Pre-Order Traversal | 0.98 (0.04) | 1.00 (0.00) | 1.00 (0.00) | ||
| Post-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Depth | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Compound | 1.00 (0.00) | 1.00 (0.00) | 0.98 (0.02) | ||
| Heap* | Compound | 0.51 (0.02) | 0.81 (0.08) | 0.66 (0.12) | |
| Heapify | 0.33 (0.00) | 0.61 (0.05) | 0.76 (0.10) | ||
| RB Tree* | Construct | 0.93 (0.00) | 0.90 (0.03) | 0.68 (0.13) | |
| Compound | 0.91 (0.04) | 0.67 (0.03) | 0.49 (0.04) | ||
| B+ Tree* | Compound | 0.97 (0.00) | 0.99 (0.02) | 0.98 (0.02) | |
| K-D Tree* | Construct | 0.97 (0.03) | 0.90 (0.00) | 0.67 (0.06) | |
| K-D Heap* | Compound | 0.23 (0.00) | 0.08 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 0.92 (0.02) | 0.98 (0.02) | 0.94 (0.02) |
| Depth-First Traversal | 0.93 (0.03) | 1.00 (0.00) | 0.96 (0.02) | ||
| DSU* | Compound | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | |
| Geom Graph* | Construct | 0.21 (0.10) | 0.08 (0.04) | 0.19 (0.13) | |
| Hybrid | Bloom Filter* | Construct | 1.00 (0.00) | 0.84 (0.07) | 0.47 (0.00) |
| DAWG* | Compound | 1.00 (0.00) | 0.03 (0.03) | 0.00 (0.00) |
A.4.2 Performance of o4-mini
@@
| Geom Graph* | Construct | 0.83 (0.07) | 0.13 (0.00) | 0.01 (0.02) | |
| Hybrid | Bloom Filter* | Compound | 1.00 (0.00) | 0.77 (0.09) | 0.07 (0.00) |
@@
| Priority Queue* | Compound | 0.89 (0.02) | 0.41 (0.02) | 0.23 (0.05) | |
| Associative | Hashmap | Compound | 0.58 (0.02) | 0.16 (0.07) | 0.11 (0.04) |
| Trie* | Compound | 0.93 (0.03) | 0.56 (0.02) | 0.37 (0.03) | |
| Suffix Tree* | Construct | 0.96 (0.04) | 0.89 (0.02) | 0.86 (0.08) | |
| Skip List* | Compound | 0.94 (0.02) | 0.87 (0.06) | 0.53 (0.03) | |
| Hierarchical | BST | Insert | 1.00 (0.00) | 0.94 (0.02) | 0.94 (0.05) |
| Remove | 0.87 (0.06) | 0.86 (0.07) | 0.89 (0.02) | ||
| In-Order Traversal | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
@@
| Post-Order Traversal | 1.00 (0.00) | 0.99 (0.02) | 0.94 (0.04) | ||
| Depth | 1.00 (0.00) | 1.00 (0.00) | 1.00 (0.00) | ||
| Compound | 0.99 (0.02) | 0.99 (0.02) | 0.98 (0.02) | ||
| Heap* | Compound | 0.78 (0.04) | 0.49 (0.10) | 0.53 (0.12) | |
| Heapify | 0.36 (0.02) | 0.06 (0.05) | 0.04 (0.04) | ||
| RB Tree* | Construct | 0.91 (0.07) | 0.53 (0.09) | 0.03 (0.03) | |
| Compound | 0.91 (0.02) | 0.41 (0.07) | 0.13 (0.03) | ||
| B+ Tree* | Compound | 1.00 (0.00) | 0.97 (0.03) | 0.94 (0.06) | |
| K-D Tree* | Construct | 0.96 (0.02) | 0.64 (0.13) | 0.16 (0.02) | |
| K-D Heap* | Compound | 0.23 (0.00) | 0.06 (0.02) | 0.00 (0.00) | |
| Network | Graph | Breadth-First Traversal | 1.00 (0.00) | 1.00 (0.00) | 0.74 (0.02) |
| Depth-First Traversal | 1.00 (0.00) | 0.81 (0.05) | 0.14 (0.02) | ||
| DSU* | Compound | 1.00 (0.00) | 0.99 (0.02) | 0.90 (0.05) | |
| Geom Graph* | Construct | 0.19 (0.02) | 0.17 (0.06) | 0.02 (0.02) | |
| Hybrid | Bloom Filter* | Construct | 0.94 (0.02) | 0.97 (0.00) | 0.66 (0.02) |
| DAWG* | Compound | 0.61 (0.05) | 0.08 (0.07) | 0.01 (0.02) |
A.4.4 Performance of Claude-3.7-Sonnet
@@
| Geom Graph* | Construct | 0.04 (0.05) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.44 (0.04) | 0.03 (0.00) | 0.00 (0.00) |
@@ A.4.5 Performance of DeepSeek-R1
| Geom Graph* | Construct | 0.99 (0.02) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.99 (0.02) | 0.92 (0.02) | 0.31 (0.02) |
@@
| Geom Graph* | Construct | 0.03 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
A.4.7 Performance of Gemini-2.0-Flash
@@
| Geom Graph* | Construct | 0.07 (0.03) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |
@@
A.4.8 Performance of Claude-3.5-Sonnet
@@
| Geom Graph* | Construct | 0.10 (0.00) | 0.00 (0.00) | 0.00 (0.00) | |
| Hybrid | Bloom Filter* | Compound | 0.10 (0.00) | 0.03 (0.00) | 0.00 (0.00) |